# Generic motives and motivic cohomology of fields

### v1

@@ 5.2. The effect of transcendental elements
We have seen in 5.2.1 that the line $n=1$ on motivic cohomology of fields has a special behavior: it does not seem to grow with transcendental elements. In general, one conjectures (after Beilinson) that for any $i>1$, $\operatorname{H}_{\mathcyr{B}}^{1,i}(K)=\operatorname{H}_{\mathcyr{B}}^{1,i}(K_{0})$ where $K_{0}\subset K$ is the field of constants (i.e. the algebraic closure of the prime subfield of $K$). This is sometimes called the rigidity conjecture. For more discussion on this and the Belinson-Soulé conjecture, the reader is advised to look at [Kah05, §4.3.4, Lem. 24]. [^41] Note that the conjectures 16 and 17 of loc. cit. are now a theorem due to Voevodsky, with important contributions of Rost.
5.2.8.
The Milnor K-theory part. The above conjecture in the restricted case $n=i$ was already asked in [BT73, Question after the proof of 5.10]. Even in that restricted case, nothing is known beyond the case $\delta(K)\leq 1$ already mentioned.
@@
Recall that the Beilinson-Soulé conjecture is equivalent to the existence of the motivic $t$-structure on (rational) Tate motives over $K$ with a weight filtration such that $\mathbbm{1}(n)$ is of weight $-2n$ (see [Lev93, Th. 4.2]). [^42] Actually, the assertion that $\mathbbm{1}(n)$ is in the heart of the motivic $t$-structure and of weight $(-2n)$ for all $n\in\mathbb{Z}$ is equivalent to the strong form of the Beilinson-Soulé conjecture for $K$.
In that respect, the novely of our question is rather about the vanishing above the Kronecker dimension of the field. Moreover, under the Beilinson-Soulé conjecture, the vanishing in this particular range is equivalent to ask that the category of Tate motives has cohomological dimension $\delta(K)$ (see again [Lev93, Cor. 4.3] in the case $\delta(K)=1$).
@@
The positive characteristic case. When $K$ is a field of characteristic $p>0$, the above conjecture is in fact a consequence of two known conjectures. And in fact, one hopes for a stronger vanishing of motivic cohomology: the conjecture of Beilinson and Parshin implies that $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)=0$ if $n\neq i$ (see [Gei98, Th. 3.4]). Therefore, the previous conjecture would be a consequence of the Beilinson-Parshin conjecture and a positive answer to the question of Bass and Tate recalled above.
@@ 5.3. Motivic cohomology with coefficients in abelian varieties
We finish this paper by mentioning an interesting cohomology theory that comes out of Voevodsky’s theory of motivic complexes.
@@
Then, given any smooth $k$-scheme of finite type, we define the motivic cohomology of $X$ with coefficients in $A$, degree $n$ and twists $i$, as the abelian group:
$H^{n,i}(X,\underline{A})=\operatorname{Hom}_{\operatorname{DM}(k)}\big{(}\operatorname{M}(X),\underline{A}(i)\big{)}.$
This new cohomology theory inherits all the properties of the motive $\operatorname{\hat{M}}(X)$:
@@
According to 4.2.4, $\underline{A}(-i)=0$ for $i>0$. One deduces from the cancellation theorem that for $i\geq 0$:
@@
where the left hand-side is the Nisnevich cohomology of the motivic complexes $\underline{A}\otimes\mathbbm{1}(i)\in\operatorname{DM^{eff}}(k)$. According to the property of the derived tensor product in $\operatorname{DM^{eff}}(k)$, one deduces that $\underline{A}(i)$ is a complex of Nisnevich sheaves concentrated in cohomological degree $\leq i$. One deduces:
@@
with the notation of 4.2.4. Moreover, it follows from 3.3.7 that $\operatorname{K}^{SM}_{*}(-;A)\simeq\widehat{\underline{A}}_{*}$ is a cycle module over $k$. One deduces a differential from Rost’s cycle complex with coefficients in the latter cycle module:
@@
Consider the above notation. One defines the codimension $n$ algebraic cycles of $X$ with coefficients in $A$ as the abelian group:
@@
The map $\mathrm{div}_{A}$ defined above will be called the divisor class map. One defines the Chow group of $X$ in codimension $n$ and coefficients in $A$ as the cokernel:
@@
Here $\sim_{A}$ is the obvious equivalence relation, that will be called the rational equivalence. As a consequence of the coniveau spectral sequence together with the previous computation, one gets:
@@
    Up to inverting the characteristic exponent of $k$, it can be extended to singular $k$-schemes in a contravariant way. The extension satisfies $\mathrm{cdh}$-descent.
@@
    One can define a compactly supported version by using motive with compact supports.
@@
The preceding definition is useful to state the computation of motivic cohomology of function fields that comes from 4.2.2.
@@
Let $K$ be a number field, and $\bar{C}/K$ be a smooth projective curve with base point $x_{0}\in\bar{C}_{(0)}$. We let $L$ be the function field of $\bar{C}$, and $A$ be the jacobian of pointed curve $(\bar{C},x_{0})$.
@@
$0\rightarrow\operatorname{H}_{\mathrm{M}}^{0,i-1}(K,\underline{A})\xrightarrow{\pi_{L/K}^{2,i}}\operatorname{H}_{\mathrm{M}}^{2,i}(L)\xrightarrow{\sum_{x}\partial_{x}}\bigoplus_{x\in C_{(0)}}\operatorname{H}_{\mathrm{M}}^{1,i-1}(\kappa_{x})\rightarrow\operatorname{H}_{\mathrm{M}}^{1,i-1}(K,\underline{A})\xrightarrow{\pi_{L/K}^{2,i}}\operatorname{H}_{\mathrm{M}}^{3,i}(L)\rightarrow 0$
@@
This is a faithful translation of the homotopy exact sequence of 4.2.2, with the duality result of 4.2.6, and using the vanishing of motivic cohomology of the fields $K$ and $\kappa_{x}$ from 5.1.4. ∎
5.3.7.
One can therefore interpret the preceding conjecture for the function field $L$ in terms of properties of the cohomology groups $\operatorname{H}_{\mathrm{M}}^{n,i}(K,\underline{A})$, which are attached to an abelian variety over a number field. These data might be easier to grasp than the transcendental extension $L$.

### v3

@@ 5.2. The effect of transcendental elements
We observed in 5.2.1 that the line $n=1$ on motivic cohomology of fields exhibits special behavior: it does not seem to grow with transcendental elements. In general, one conjectures following Beilinson that for any $i>1$, $\operatorname{H}_{\mathcyr{B}}^{1,i}(K)=\operatorname{H}_{\mathcyr{B}}^{1,i}(K_{0})$ where $K_{0}\subset K$ is the field of constants (i.e. the algebraic closure of the prime subfield of $K$). This is sometimes called the rigidity conjecture. For more discussion on this and the Belinson-Soulé conjecture, the reader is referred to [Kah05, §4.3.4, Lem. 24]. [^50] Note that the conjectures 16 and 17 of loc. cit. are now established as a theorem by Voevodsky, with significant contributions from Rost.
@@
The Milnor K-theory part. The above conjecture in the special case $n=i$ was already formulated in [BT73, Question after the proof of 5.10]. Even in this special case, nothing is known beyond the case $\delta(K)\leq 1$ already mentioned.
@@
Recall that the Beilinson-Soulé conjecture is equivalent to the existence of a motivic $t$-structure on (rational) Tate motives over $K$ with a weight filtration such that $\mathbbm{1}(n)$ is of weight $-2n$ (see [Lev93, Th. 4.2]). [^51] In fact, the assertion that $\mathbbm{1}(n)$ lies in the heart of the motivic $t$-structure and has weight $(-2n)$ for all $n\in\mathbb{Z}$ is equivalent to the strong form of the Beilinson-Soulé conjecture for $K$.
In this respect, the novelty of our question concerns the vanishing above the Kronecker dimension of the field. Moreover, under the Beilinson-Soulé conjecture, the vanishing in this particular range is equivalent to the assertion that the category of Tate motives has cohomological dimension $\delta(K)$ (see again [Lev93, Cor. 4.3] in the case $\delta(K)=1$).
5.2.11.
The positive characteristic case. When $K$ is a field of characteristic $p>0$, the above conjecture is actually a consequence of two well-known conjectures. In fact, a stronger vanishing of motivic cohomology is anticipated. The Beilinson-Parshin conjecture implies that $\operatorname{H}_{\mathcyr{B}}^{n,i}(K)=0$ if $n\neq i$ (see [Gei98, Th. 3.4]). Thus, the previous conjecture would follow from the Beilinson-Parshin conjecture and a positive resolution of the question of Bass and Tate mentioned above.
@@ 5.3. Motivic cohomology with coefficients in abelian varieties
We finish this paper by mentioning an interesting cohomology theory that comes from Voevodsky’s theory of motivic complexes.
@@
Then, for any smooth $k$-scheme of finite type, we define the motivic cohomology of $X$ with coefficients in $A$, degree $n$ and twist $i$, as the abelian group:
$H^{n,i}(X,\underline{A})=\operatorname{Hom}_{\operatorname{DM}(k)}\big{(}\operatorname{M}(X),\underline{A}(i)[n]\big{)}.$
This new cohomology theory inherits all the properties of the motive $\operatorname{M}(X)$:
@@
According to 4.3.4, $\underline{A}(-i)=0$ for $i>0$. In addition, one deduces from the cancellation theorem that for $i\geq 0$:
@@
where the left-hand side is the Nisnevich cohomology of the motivic complexes $\underline{A}\otimes\mathbbm{1}(i)\in\operatorname{DM^{eff}}(k)$. According to the property of the derived tensor product in $\operatorname{DM^{eff}}(k)$, one deduces that $\underline{A}(i)$ is a complex of Nisnevich sheaves concentrated in cohomological degree $\leq i$. One deduces:
@@
with the notation of 4.3.4. Moreover, it follows from 3.3.6 that $\operatorname{K}^{SM}_{*}(-;A)\simeq\widehat{\underline{A}}_{*}$ is a cycle module over $k$. One deduces a differential from Rost’s cycle complex with coefficients in the latter cycle module:
@@
Consider the above notation. One defines the $n$-codimensional algebraic cycles of $X$ with coefficients in $A$ as the abelian group:
@@
The map $\mathrm{div}_{A}$ defined above is called the divisor class map. One defines the Chow group of $X$ in codimension $n$ and coefficients in $A$ as the cokernel:
@@
Here $\sim_{A}$ is the obvious equivalence relation, that we call rational equivalence. As a consequence of existence of the coniveau spectral sequence together with the previous computation, one gets:
@@
    Up to inverting the characteristic exponent of $k$, the preceding definition can be extended to singular $k$-schemes in a contravariant way. The corresponding extended cohomology theory satisfies $\mathrm{cdh}$-descent.
@@
    One can define a compactly supported version by using motives with compact support.
@@
The preceding definition is useful to state the computation of motivic cohomology of function fields given in 4.3.2.
@@
Let $K$ be a number field, and $\bar{C}/K$ be a smooth projective curve with a base point $x_{0}\in\bar{C}_{(0)}$. We let $L$ be the function field of $\bar{C}$, and $A$ be the jacobian of the pointed curve $(\bar{C},x_{0})$.
@@
$0\rightarrow\operatorname{H}_{\mathrm{M}}^{0,i-1}(K,\underline{A})\xrightarrow{\pi_{L/K}^{2,i}}\operatorname{H}_{\mathrm{M}}^{2,i}(L)\xrightarrow{\sum_{x}\partial_{x}}\bigoplus_{x\in C_{(0)}}\operatorname{H}_{\mathrm{M}}^{1,i-1}(\kappa_{x})\rightarrow\operatorname{H}_{\mathrm{M}}^{1,i-1}(K,\underline{A})\xrightarrow{\pi_{L/K}^{3,i}}\operatorname{H}_{\mathrm{M}}^{3,i}(L)\rightarrow 0$
@@
This follows directly from the homotopy exact sequence of 4.3.2, with the duality result of 4.3.7, and using the vanishing of motivic cohomology of the fields $K$ and $\kappa_{x}$ from 5.1.4. ∎
Remark 5.3.7.
One can therefore interpret the 5.2.6 for the function field $L$ in terms of properties of the cohomology groups $\operatorname{H}_{\mathrm{M}}^{n,i}(K,\underline{A})$, which are attached to an abelian variety over a number field. These cohomology groups may be more accessible.